# On-board Satellite Image Classification for Earth Observation: A Comparative Study of Pre-trained Vision-Transformer Models

### v1

@@ Abstract
Remote sensing image classification is a critical component of Earth observation (EO) systems, traditionally dominated by convolutional neural networks (CNNs) and other deep learning (DL) techniques. However, the advent of Transformer-based architectures and large-scale pre-trained models has significantly shifted, offering enhanced performance and efficiency. This study focuses on identifying the most effective pre-trained model for land use classification for onboard-satellite (On-Air) processing, emphasizing achieving high accuracy, computational efficiency, and robustness against noisy data—conditions commonly encountered during On-Air inference. Through extensive experimentation, we compared traditional CNN-based models, ResNet-based models, and various pre-trained vision Transformer models. Our findings demonstrate that pre-trained Transformer models, particularly MobileViTV2 and EfficientViT-M2, outperform models trained from scratch in accuracy and efficiency. These models achieve high performance with reduced computational requirements and exhibit greater resilience during inference under noisy conditions. While MobileViTV2 excelled on clean validation data, EfficientViT-M2 proved more robust when handling noise, making it the most suitable model for EO tasks. In conclusion, EfficientViT-M2 is the optimal choice for reliable and efficient On-Air remote-sensing image classification, achieving 98.76% accuracy, precision, and recall. Specifically, EfficientViT-M2 delivered the highest performance across all metrics, excelled in training efficiency (1,000s) and inference time (10s), and demonstrated greater robustness (overall robustness score at 0.79).
@@
EO, remote sensing, Transformers, on-board processing, pre-trained ViT, model robustness
@@ I Introduction
Remote sensing image classification (RSIC) has long been a critical application in EO systems, with traditional methods primarily relying on CNNs and other DL techniques. These approaches have been well-documented and validated over the years, demonstrating significant capabilities in processing and classifying remote sensing (RS) data with reasonable accuracy and efficiency [1, 2, 3]. However, as the field has evolved, so too have the methods, with recent advances emphasizing the use of Transformer-based architectures, which have begun to surpass traditional CNNs in many image classification (IC) tasks, including RS [4, 5].
@@
The rise of large-scale pre-trained models has marked a new era in artificial intelligence (AI), achieving breakthroughs across various domains, particularly in natural language processing (NLP) with models like BERT [6] and GPT [7]. This paradigm shift is now extending into computer vision, where pre-trained vision Transformer (ViT) models are proving to be highly effective [8, 9]. These models benefit from extensive pre-training on large datasets, allowing them to learn rich contextual representations that can be fine-tuned for specific tasks with minimal additional training. The success of these models is primarily attributed to advancements in computational power, the availability of vast amounts of data, and innovations in model architecture and efficiency [10].
In RS, while pre-trained models, especially pre-trained ViT models, hold great potential for enhancing classification performance and reducing computational demands during inference, their application on board satellites is constrained due to the significant power consumption required. Implementing such large models onboard is particularly challenging, as seen with the pioneering use of neural networks in satellite-based EO systems. For instance, the $\Phi$-Sat-1 mission employed the CloudScout model, a CNN-based neural network, for image segmentation, marking the first instance of On-Air DL [11]. Similarly, the $\Phi$-Sat-2 mission utilized a convolutional autoencoder for image compression, further illustrating the practical challenges of deploying computationally intensive models in space environments [12]. One helpful method to enhance computational efficiency in such downstream tasks is transfer learning, where a pre-trained model is fine-tuned for a specific application, such as land use classification. This approach has shown consistent improvements in performance across various tasks, as it leverages the pre-learned features from large-scale datasets, requiring only minimal adaptation to the new task [13].
Figure 1: EO system architecture in which the recorded EO data is preprocessed and inference at the satellite (Fig from Prabhu). Redrwaing the Figure, removing the level of pre-processing
@@
This study aims to identify the most effective pre-trained model for RSIC, explicitly focusing on land use classification applications with On-Air processing. The goal is to find a model that provides high accuracy, operates efficiently with limited computational resources, and demonstrates robustness under noisy conditions, which is common in satellite-based inference due to environmental factors or instrument imperfections. As shown in Fig. 1, the process begins with the experimental training of the model on the ground, where various pre-trained models are rigorously evaluated. Once the most effective model is identified, it is deployed to On-Air inference. In this setup, the EO satellite records sensor data, which is then preprocessed by the onboard processor, eliminating duplicates and preparing the data for level 0/1 processing. The preprocessed data is subsequently transmitted back to Earth, where further processing and analysis are conducted, ultimately supporting various user applications. By leveraging the trained model, we aim to optimize the classification performance directly within the satellite, ensuring efficient and accurate land use classification in real-time.
@@
To achieve this, we conducted extensive experiments comparing traditional CNN-based, ResNet-based, and Transformer-based models for IC. The study then delves into various pre-trained ViT models to identify the optimal balance between model complexity, computational efficiency, and performance. A key aspect of our evaluation is the robustness of these models when faced with noisy data during inference, as this scenario closely mirrors the real-world conditions of satellite EO systems.
@@
Our findings indicate that pre-trained vision Transformer models significantly outperform models trained from scratch, particularly in efficiency and accuracy. Among the models tested, MobileViTV2 and EfficientViT-M2 emerged as the top performers for On-Air EO processing. These models deliver high performance with lower computational demands and exhibit greater robustness in the presence of noisy test data. While MobileViTV2 achieved higher validation performance on clean data, EfficientViT-M2 proved more resilient during inference with noisy data, making it the most suitable model for our specific application.
@@
In conclusion, EfficientViT-M2 is identified as the optimal model for On-Air EO processing tasks due to its balance of training efficiency, high accuracy, and robustness under challenging inference conditions. This model offers a practical solution for enhancing the reliability and effectiveness of RSIC in real-world satellite operations.
@@ II Related Works
Recent RS advancements have focused on enhancing CNN models through transfer learning, structural modifications, and integrating attention mechanisms. For instance, [14] proposed a self-attention fused CNN architecture optimized explicitly for land use and land cover classification, incorporating advanced data augmentation and custom-designed CNN models to improve classification performance. Similarly, [15] introduced a CNN architecture that combines dense blocks with inverted bottleneck residuals, further enhancing deep feature extraction for RS data.
@@
One of the critical benefits of transfer learning is its ability to mitigate the domain gap between natural and RS images. This gap has been a significant challenge due to the lack of large-scale, widely recognized benchmarks like ImageNet within the RSI community. To address this, [16] proposed the MGC method, which employs an multilayer perceptron (MLP) to guide the pre-training of a CNN using small-scale RS datasets. By leveraging attention guidance, MGC effectively directs CNN branches to focus on foreground regions, thereby learning more discriminative representations.
Another notable example is the transfer learning approach presented by [17], which utilizes fully pre-trained deep convolutional networks for land-use classification in high spatial resolution imagery. This method addresses the common issue of separation between the feature descriptor and classifier parts in transferred CNNs by pre-training both components, resulting in faster convergence and improved accuracy without sacrificing performance.
@@
Despite the success of CNNs, the inherent limitations of these models, particularly when initialized with ImageNet pretraining, highlight the need for more specialized approaches in RS. Domain gaps between RS and natural images often lead to suboptimal performance when using IMP, prompting researchers to explore alternative transfer learning strategies. In this context, MLP models have shown promise in addressing challenges such as cross-domain few-shot classification. For example, [18] demonstrated that MLPs could significantly enhance discriminative capabilities and alleviate distribution shifts, thereby improving classification performance in RS tasks. Integrating ViT models and advanced transfer learning techniques has become increasingly popular as the field progresses. For instance, [19] trained a model from scratch using a Transformer-based architecture combined with a CNN classifier layer, demonstrating the effectiveness of such hybrid approaches in processing EuroSAT images.
Some studies have focused on generating synthetic data or integrating multiple data sources to maximize data utility further. For example, [20] explored the scaling laws of synthetic images generated by state-of-the-art text-to-image models, highlighting the potential of synthetic data in scenarios with limited authentic images. Similarly, [21] introduced the GeRSP framework, which combines self-supervised and supervised pretraining branches to learn robust representations from both RS and natural images, enhancing model initialization for various downstream tasks.
@@
Moreover, addressing the challenges of data scarcity and computational efficiency has led to the development of innovative models. [22] proposed the GeoSystemNet model, which applies DL in the context of geosystem analysis to overcome labeled data scarcity in high-resolution RS data classification. In parallel, [23] introduced a parameter-efficient continual learning framework that dynamically expands pre-trained models, such as CLIP, with Mixture-of-Experts adapters to handle new tasks while preserving zero-shot recognition capabilities. In addition to improving model performance, recent research has emphasized the importance of interpretability in RS models. [24] developed an interpretable DL framework for land use classification using SHAP (Shapley Additive Explanations), offering insights into how different spectral bands affect model predictions.
While CNN-based approaches continue to dominate RSIC, incorporating Vision Transformers, advanced transfer learning methods, and synthetic data generation represents a significant evolution in the field. These developments enhance model robustness and generalizability and pave the way for more efficient and reliable RS applications.
@@
Given the rapid advancements in ViT models and their increasing application in RS, there is a critical need for comprehensive studies that extensively compare existing ViT models, particularly in onboard processing. Current trends emphasize model complexity reduction through downscaling, making these models more suitable for resource-constrained environments such as satellites. However, despite these efforts, there remains a significant gap in understanding how these optimized models perform under real-world conditions, where noisy data is a common challenge. Therefore, it is essential to evaluate these models based on their computational efficiency and rigorously test their robustness during inference with noisy data. Such an analysis will provide valuable insights into the practical deployment of ViT models for On-Air EO processing, ensuring high performance and reliability in challenging operational environments.
@@ III-A Dataset
Public Benchmark for Land Use and Land Cover Classification Dataset [25]. The EuroSAT dataset is a large-scale benchmark dataset designed explicitly for land use and land cover classification derived from Sentinel-2 satellite imagery. It comprises 27,000 geo-referenced labeled images, each measuring 64x64 pixels and spanning 13 spectral bands. The dataset is categorized into 10 classes. Each class contains 2000–3000 images, including industrial buildings, residential buildings, annual crops, permanent crops, rivers, seas and lakes, Herbaceous vegetation, highways, pastures, and forests across Europe. Due to its compact image size and diverse class representation, EuroSAT is particularly well-suited for developing and evaluating DL models intended for onboard-processing EO missions. This makes it a valuable resource for applications such as real-time environmental monitoring, disaster response, and precision agriculture, where in-situ processing capabilities are crucial for timely decision-making.
@@ III-B Machine Learning Models
In this study, we utilize a diverse set of machine learning (ML) models, including both models trained from scratch and pre-trained models, to evaluate their performance on IC tasks. The models range from traditional convolutional neural networks (CNNs) to advanced architectures like Vision Transformers (ViTs) and hybrid models that combine convolutional layers with transformer-based processing. This selection allows us to explore the effectiveness of different architectures, from lightweight and efficient models to more complex, high-capacity networks, in handling IC challenges.

### v2

# Onboard Satellite Image Classification for Earth Observation: A Comparative Study of ViT Models

@@ Abstract
Remote sensing (RS) image classification (IC) is a critical component of Earth observation (EO) systems, traditionally dominated by convolutional neural networks (CNNs) and other deep learning (DL) techniques. However, the advent of Transformer-based architectures and large-scale pre-trained models has significantly shifted the trend by offering enhanced performance and efficiency. Hence, this study focuses on identifying the most effective pre-trained model for land use classification in onboard satellite processing, emphasizing achieving high accuracy, computational efficiency, and robustness against noisy data—conditions commonly encountered during satellite-based inference. Through extensive experimentation, we compare the performance of traditional CNN-based, ResNet-based, and various pre-trained vision Transformer models. Our findings demonstrate that pre-trained Vision Transformer (ViT) models, particularly MobileViTV2 and EfficientViT-M2, outperform models trained from scratch in terms of accuracy and efficiency. These models achieve high performance with reduced computational requirements and exhibit greater resilience during inference under noisy conditions. While MobileViTV2 has excelled on clean validation data, EfficientViT-M2 has proved to be more robust when handling noise, making it the most suitable model for onboard satellite EO tasks. Our experimental results demonstrate that EfficientViT-M2 is the optimal choice for reliable and efficient RS-IC in satellite operations, achieving 98.76% of accuracy, precision, and recall. Precisely, EfficientViT-M2 delivers the highest performance across all metrics, excels in training efficiency (1,000s) and inference time (10s), and demonstrates greater robustness (overall robustness score of 0.79). Consequently, EfficientViT-M2 consumes 63.93% less power than MobileViTV2 (79.23 W) and 73.26% less power than SwinTransformer (108.90 W). This highlights its significant advantage in energy efficiency.
@@
Earth observation, remote sensing, Transformers, on-board processing, pre-trained ViT, model robustness
@@ I Introduction
The increasing demand for satellites to support EO and RS missions has driven a rapid expansion in the number of Low Earth Orbit (LEO) satellites [1]. These missions are critical for various applications, including environmental monitoring, disaster response, precision agriculture, and scientific research, where high-frequency, high-resolution data is essential. A key component of EO systems is RS-IC, which plays a fundamental role in analyzing satellite-acquired data. Traditionally, IC in RS has relied on CNNs and other DL techniques, which have been widely validated and proven effective for processing and classifying RS data with high accuracy and efficiency [2, 3, 4].
@@
However, recent advances have shifted towards Transformer-based architectures, surpassing conventional CNNs in several IC tasks, including those in RS applications [5, 6]. The rise of large-scale pre-trained models has marked a new era in artificial intelligence (AI), achieving breakthroughs across various domains, particularly in natural language processing with models like BERT [7]. This paradigm shift extends into computer vision, where pre-trained ViTsize models demonstrate superior performance across a wide range of tasks [8, 9]. These models benefit from extensive pre-training on large datasets, allowing them to learn rich contextual representations that can be fine-tuned for specific tasks with minimal additional training. The success of these models is primarily attributed to advancements in computational power, the availability of vast amounts of data, and innovations in model architecture and efficiency [10]. While pre-trained ViT models hold great potential for enhancing classification performance and reducing computational demands during inference, their application in onboard satellite systems (OSS) is constrained due to such large-scale models’ high power consumption and computing resource requirements. Implementing these models for OSS real-time decision-making remains a significant challenge.
Furthermore, a major operational bottleneck for LEO satellites is their dependence on ground stations for data transmission, leading to limited communication opportunities. The short visibility window results in prolonged connectivity loss, delaying access to critical data and hindering timely responses [11].
Satellite Internet Providers (SIPs) emergence, e.g. SpaceX and OneWeb, offers the potential to overcome these limitations by enabling continuous (24/7) connectivity, allowing on-demand data access [12]. However, seamless connectivity alone is insufficient to meet the modern EO and RS requirements, which demand low-latency, real-time decision-making capabilities. This highlights the need for advanced onboard processing solutions that can autonomously analyze data, prioritize critical information, and take immediate action, such as refocusing during the subsequent revisit, without relying on delayed ground-based interventions [13].
Therefore, most onboard neural networks (NNs) are designed to be lightweight, prioritizing computational efficiency over model complexity. For example, the $\Phi$-Sat-1 mission employed the CloudScout model, a CNN-based NN, for onboard image segmentation on the Intel Movidius Myriad 2 vision processing unit (VPU), making it the first instance of DL deployed on a satellite [14]. Similarly, the $\Phi$-Sat-2 mission implemented a convolutional autoencoder for image compression to minimize data transmission needs [15] and tested on three different hardware, including graphic processing unit (GPU) NVIDIA GeForce GTX 1650 VPU Myriad 2, and central processing unit (CPU) Intel Core i7-6700. These examples highlight the practical limitations in deploying more complex models, such as Transformer-based architectures and onboard satellites, due to restricted power budgets and limited processing resources. Consequently, while CNN-based models can achieve reasonable performance, they cannot often capture the intricate patterns and contextual dependencies required for more advanced tasks, limiting their effectiveness in real-world remote sensing applications. One promising approach to onboard processing is using neuromorphic processors, which rely on converting traditional NNs into Spiking NNs (SNNs). Currently, this conversion is best supported for CNN-based models, as demonstrated in recent studies on energy-efficient neuromorphic computing for space applications [16, 17]. While these lightweight CNN-based models are advantageous due to their low power consumption and reduced computational requirements, they often fall short in terms of prediction accuracy and robustness compared to more complex Transformer-based models, which are better suited for capturing intricate patterns and contextual dependencies in RS data.
One promising approach to enhance computational efficiency for such downstream tasks is transfer learning, where a pre-trained model is fine-tuned for a specific application. This method leverages the rich, hierarchical features learned from large-scale datasets, minimizing the need for extensive retraining and enabling rapid adaptation to new tasks with significantly fewer labeled samples [18]. Transfer learning has consistently demonstrated improvements in performance across various tasks, making it particularly beneficial for resource-constrained environments like OSS, where both data and computational power are limited.
@@
Building on this strategy, this study aims to identify the most effective pre-trained model for RS-IC, explicitly focusing on land-use classification applications in OSS. The goal is to find a model that provides high accuracy, operates efficiently with limited computational resources, and demonstrates robustness under noisy conditions, which is common in satellite-based inference due to environmental factors or hardware imperfections. The process begins with the experimental training of the model on the ground, where various pre-trained models are rigorously evaluated. Once the most effective model is identified, it is deployed in the onboard satellite processor for real-time inference. In this setup, the EO satellite records sensor data, which the onboard processor then preprocesses. By leveraging the trained model on the ground, we aim to optimize the inference’s classification performance and power consumption efficiency directly within the satellite processor, ensuring efficient and accurate real-time RS-IC.
@@
To achieve this, we conduct extensive experiments comparing traditional CNN-based, ResNet-based, and Transformer-based models for IC. The study then delves into various pre-trained ViT models to identify the optimal balance between model computational efficiency and performance. A key aspect of our evaluation is the robustness of these models when faced with noisy data during on-board inference.
@@
Our findings indicate that pre-trained ViT models significantly outperform models trained from scratch, particularly in classification accuracy, precision, and recall. Among the models tested, MobileViTV2 [19] and EfficientViT-M2 [20] have emerged as the top performers for onboard satellite EO. These models deliver high performance with lower computational demands and exhibit greater robustness in the presence of noisy test data. While MobileViTV2 has achieved higher validation performance on clean data, EfficientViT-M2 has proved more resilient during inference with noisy data, making it the most suitable model for our specific application.
@@
In conclusion, EfficientViT-M2 is identified as the optimal model for EO-OSSs due to its balance of training efficiency, high accuracy, and robustness under challenging inference conditions. This model offers a practical solution for enhancing the reliability and effectiveness of RS-IC in real-world satellite operations. We summarize our contributions in terms of translational AI for onboard satellite processing as follows:
  • Adaptation of pre-trained models for resource-constrained environments: We demonstrate how large-scale, pre-trained Transformer models can be effectively adapted for onboard satellite processing, providing a new framework for deploying complex AI models in power- and computation-limited environments while maintaining high accuracy and efficiency.
  • Optimizing real-time decision-making in satellite operations: EfficientViT-M2 is identified as the optimal model for onboard inference, demonstrating robustness with noisy sensor data and enabling autonomous, real-time decisions. In scenarios like forest fires or oil spills, the satellite can autonomously repoint its sensors or switch to infrared imaging for better visibility without waiting for ground-based instructions, significantly improving disaster response and environmental monitoring efficiency.
  • Bridging the gap between ground and onboard processing: By conducting rigorous ground-based evaluations of state-of-the-art models and then translating these findings to onboard implementation, we bridge the gap between theoretical AI research and practical applications in space, highlighting a pathway for deploying advanced AI solutions to enhance operational efficiency and autonomy in satellite EO missions.
The remainder of this paper is organized as follows. Section II provides a brief survey of related works, serving as the basis for highlighting this paper’s contributions. Section III presents the methodologies, including the dataset, machine learning (ML) models, and evaluation metrics employed in this comprehensive study. The experimental setup is explained in Section IV, followed by the numerical results and discussion in Section V. Finally, Section VI concludes the paper. Reproducible codes for data augmentation, noisy data, training, and inference are available from our shared Github repository 111 https://github.com/ltdung/SnT-SENTRY.
@@ II Related Works
Recent advancements in RS have focused on enhancing CNN models through transfer learning, structural modifications, and integrating attention mechanisms. For instance, [21] proposed a self-attention fused CNN architecture optimized explicitly for hyperspectral land use and land cover classification, incorporating advanced data augmentation and custom-designed CNN models to improve classification performance. Similarly, [22] introduced a CNN architecture that combines dense blocks with inverted bottleneck residuals, further enhancing deep feature extraction for RS data.
@@
One of the critical benefits of transfer learning is its ability to mitigate the domain gap between natural and RS images. This gap has been a significant challenge due to the lack of large-scale, widely recognized benchmarks like ImageNet within the RS community. To address this, [23] proposes the MGC method, which employs a multilayer perceptron (MLP) to guide the pre-training of a CNN using small-scale RS datasets. By leveraging attention guidance, MGC effectively directs CNN branches to focus on foreground regions, thereby learning more discriminative representations. Another notable example is the transfer learning approach presented by [24], which utilizes fully pre-trained deep CNNs for land-use classification in high spatial resolution imagery. This method addresses the common issue of separation between the feature descriptor and classifier parts in transferred CNNs by pre-training both components, resulting in faster convergence and improved accuracy without sacrificing performance.
@@
Despite the success of CNNs, the inherent limitations of these models, particularly when initialized with ImageNet pretraining, highlight the need for more specialized approaches in RS. Domain gaps between RS and natural images often lead to suboptimal performance when using ImageNet pretraining, prompting researchers to explore alternative transfer learning strategies. In this context, MLP models have shown promise in addressing challenges such as cross-domain few-shot classification. For example, [25] demonstrated that MLPs could significantly enhance discriminative capabilities and alleviate distribution shifts, thereby improving classification performance in RS tasks. Furthermore, integrating ViT models and advanced transfer learning techniques has become increasingly popular as the field progresses. For instance, [26] trained a model from scratch using a Transformer-based architecture combined with a CNN classifier layer, demonstrating the effectiveness of such hybrid approaches in processing EuroSAT images. Additionally, some studies have focused on generating synthetic data or integrating multiple data sources to maximize data utility further. For example, [27] explored the scaling laws of synthetic images generated by state-of-the-art text-to-image models, highlighting the potential of synthetic data in scenarios with limited authentic images. Similarly, [28] introduced the GeRSP framework, which combines self-supervised and supervised pretraining branches to learn robust representations from both RS and natural images, enhancing model initialization for various downstream tasks.
@@
Moreover, addressing the challenges of data scarcity and computational efficiency has led to the development of innovative models. [29] proposes the GeoSystemNet model, which applies DL in the context of geosystem analysis to overcome labeled data scarcity in high-resolution RS data classification. In parallel, [30] introduces a parameter-efficient continual learning framework that dynamically expands pre-trained models, such as CLIP, with Mixture-of-Experts adapters to handle new tasks while preserving zero-shot recognition capabilities. In addition to improving model performance with computational efficiency, recent research has emphasized the importance of interpretability in RS models. [31] develops an interpretable DL framework for land use classification using SHAP (Shapley Additive Explanations), offering insights into how different spectral bands affect model predictions. While CNN-based approaches continue to dominate RS-IC, incorporating ViTs, advanced transfer learning methods, and synthetic data represents a significant evolution in the field. These developments enhance model robustness and generalizability and pave the way for more efficient and reliable RS applications.
@@
Given the rapid advancements in ViT models and their increasing application in RS, there is a critical need for comprehensive studies that extensively compare existing ViT models, particularly in onboard processing. Current trends emphasize model complexity reduction through downscaling, making these models more suitable for resource-constrained environments such as satellites. However, despite these efforts, there remains a significant gap in understanding how these optimized models perform under real-world conditions, where noisy data is a common challenge. Therefore, it is essential to evaluate these models based on their computational efficiency and rigorously test their robustness during inference with noisy data. Such an analysis will provide valuable insights into the practical deployment of ViT models for onboard satellite EO, ensuring high performance and reliability in challenging operational environments.
@@ III Methodology
This section details the methodological framework used in evaluating pre-trained models for land use classification in onboard satellite processing. We begin by outlining the dataset and preprocessing steps, followed by a list of the ML models that can be utilized for RS-IC. Additionally, we present the evaluation metrics employed to assess model performance in terms of accuracy, computational efficiency, and robustness, particularly under noisy conditions commonly encountered in satellite-based inference.
@@ III-A Dataset
The public benchmark for land use and land cover classification dataset, EuroSAT [32], will be used in this study. The EuroSAT dataset is a large-scale benchmark dataset designed explicitly for land use and land cover classification derived from Sentinel-2 satellite imagery. It comprises 27,000 geo-referenced labeled images, each measuring 64x64 pixels and spanning 13 spectral bands. The dataset is categorized into 10 classes. Each class contains 2000–3000 images, including industrial buildings, residential buildings, annual crops, permanent crops, rivers, seas and lakes, Herbaceous vegetation, highways, pastures, and forests across Europe. Due to its compact image size and diverse class representation, EuroSAT is particularly well-suited for developing and evaluating DL models intended for onboard satellite processing in EO missions. This makes it a valuable resource for applications such as real-time environmental monitoring, disaster response, and precision agriculture, where in-situ processing capabilities are crucial for timely decision-making.
@@ III-B Machine Learning Models
CNNs have long been the benchmark for visual data processing. However, recent studies demonstrate that ViTs can achieve comparable or even superior performance in image classification tasks [33]. In this study, we evaluate a diverse set of ML models, including both models trained from scratch and pre-trained models, across various IC tasks. The selection spans traditional CNNs to advanced architectures like ViTs and hybrid models combining convolutional layers with Transformer-based processing. This allows us to assess the effectiveness of lightweight, efficient models alongside more complex, high-capacity networks in addressing IC challenges.